A ship captain is plotting a course to an island 80 km SW of his current position.

The wind is blowing from the East at 15 km/h. What must the velocity vector of

the ship be in order for it to arrive at the island in 2 hours (hint: think about what

the resultant velocity must be).


1
Expert's answer
2022-05-03T03:00:01-0400

Resultant velocity is "|v|=\sqrt{|a|^2+|b|^2-2|a||b|cos \alpha}"

"|v|=80\/2=40 km\/h"

"40=\sqrt{|a|^2+15^2-2|a|x15cos (45)}"

"1600=|a|^2+225-21.3|a|"

"|a|^2-21.3|a|-1375=0"

"D=21.3^2+4x1375=5953.69"

"|a|_1=\frac{21.3+\sqrt{5953.69}}{2}=49.23"

"|a|_2=\frac{21.3-\sqrt{5953.69}}{2}=-27.94<0"

Answer: velocity vector of the ship should be 49.23 km/h to SW.